OpenStudy (anonymous):

What is a valid statement of the Pythagorean Theorem? A.In a right triangle, where a and b are the lengths of the legs and c is the length of the hypotenuse, √c = √a + √b. B.In a right triangle, where a and b are the lengths of the legs and c is the length of the hypotenuse, c2= a2 + b2. C.In a right triangle, where a and b are the lengths of the legs and c is the length of the hypotenuse, √c = a2 + b2. D.The distance between two points is the square of the difference of their x and y coordinates.

OpenStudy (lgbasallote):

the pythagorean theorem is \[a^2 + b^2 = c^2\] does that help?

OpenStudy (anonymous):

yes but idk what the answer could be i know the pythagorean theorem already

OpenStudy (lgbasallote):

well which one of your choices says \[a^2 + b^2 = c^2\]

OpenStudy (anonymous):

A but, the question says which one is valid

OpenStudy (lgbasallote):

nope. A says \[\sqrt c= \sqrt a + \sqrt b\] pythagorean theorem is \[c^2 = a^2 + b^2\] they're different

OpenStudy (anonymous):

would it be C